OpenStudy (anonymous):

A college bought some Apple Ipads(345$ each), twice and many Kindles(150$ each) and half as many Nooks(150$ each) as Ipads. If the total was 15,600$, how many of each reader was purchased?

OpenStudy (helder_edwin):

let \(x\) be the number of Ipads, then the college bought \(2x\) kindles and \(\frac{x}{2}\) nooks. then \[ \large 345x+150(2x)+150\frac{x}{2}=15600 \]

OpenStudy (anonymous):

I added all those up and got 720x=15600 are you sure thats correct?

OpenStudy (anonymous):

his question looks right, just divide the right side of the equation by 720.

OpenStudy (anonymous):

A = 2K A = N/2 -> N = 2A = 4K 345A + 150K + 150N = 15,600 345 ( 2k) + 150 K + 150 ( 4K) = 15,600